# Prediction Sets and Conformal Inference with

### v3

@@ 4 Conformal inference with valued outcomes
where $\operatorname{P}$ denotes the joint probability of $\left\{\bar{Z}_{i}=(X_{i},Y_{i},Y^{L}_{i},Y_{i}^{U}):i\in\mathcal{I}\cup\left\{{n+1}\right\}\right\}$. Notice that this is the finite-sample validity condition typically considered in conformal inference literature where the probability is the joint distribution for both the observed sample and the individual to be predicted. While the oracle prediction set $C^{*}(x)$ depends on the true distribution $P$, conformal inference constructions lead to a prediction set that is finite-sample valid for all $P$, making it distribution-free.
In particular, we will achieve finite-sample validity using the method of split conformal inference (see Romano et al., (2019) and Lei et al., (2018)). One feasible split conformal inference procedure for the oracle prediction interval $C^{*}_{I}$ in our setting can be described as follows. We first split the sample $\mathcal{D}$ into a training set $\mathcal{D}_{1}=\left\{Z_{i}:i\in\mathcal{I}_{1}\right\}$ and a calibration set $\mathcal{D}_{2}=\left\{Z_{i}:i\in\mathcal{I}_{2}\right\}$. We calculate the estimator $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$ for the oracle prediction set with the training set for $x\in\mathcal{X}$ as in Equation 7. The following scores $s_{j}=s(Y^{L}_{j},Y^{U}_{j},X_{j};\hat{\tau}_{0},\hat{\tau}_{1})$ are then computed for each $j\in\mathcal{I}_{2}$ in the calibration set, given $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$, where
@@
This score function is a modified version of the score function used in the case of conformalized quantile regression with no interval censoring studied in Romano et al., (2019). The score function is negative when the interval $[y^{L},y^{U}]$ is contained in $[t_{0}(x),t_{1}(x)]$ and positive otherwise. By allowing for negative scores, the conformal prediction set $\tilde{C}_{I}(x)$ penalises both overcoverage and undercoverage.
@@
$\vartheta_{1-\alpha}=q\left(\left\{s_{j}:j\in\mathcal{I}_{2}\right\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert)\right),$
@@
$s(y^{L},y^{U},x;C)=\min_{m\leq M}\max\left\{t_{0m}(x)-y^{L},y^{U}-t_{1m}(x)\right\}.$ (12)
Notice that $s_{j}\leq 0$ if and only if $[Y^{L}_{j},Y^{U}_{j}]\subset[t_{0m},t_{1m}]$ for some $m\leq M$. We can then compute $\vartheta_{1-\alpha}=q(\{s_{j}:j\in\mathcal{I}_{2}\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert))$. The conformal prediction set is then given by  [^2] Here, the endpoints of the estimated $\hat{C}$ are adjusted uniformly using the same $\vartheta_{1-\alpha}$. A more sophisticated adjustment method is discussed in Appendix C, where we propose a minimal Euclidean distance adjustment. We conjecture that this approach could improve the convergence rate of the conformal prediction set. The optimal choice of conformity score remains an interesting question.
@@
Under the assumption that $\{(X_{i},Y_{i},Y_{i}^{L},Y_{i}^{U}):i\in\mathcal{I}\cup\{n+1\}\}$ are i.i.d. [^3] In order to perform conformal inference, it is sufficient that the sample satisfies the weaker condition of exchangeability. Here independence is assumed for simplicity and for establishing the asymptotics for the nonparametric estimators we will use. Barber et al., (2023) studies conformal inference beyond exchangeability and is applicable for time series data. Our results can be readily extended to the time series settings combined with their procedure. , the conformal prediction set $\tilde{C}$ in Equation 13 is finite-sample marginally valid, that is, for all $P$,
@@
The following theorem shows that the conformal prediction set $\tilde{C}(x)$ preserves the consistency of the estimator $\hat{C}(x)$ for the oracle prediction set. Recall the definition of $s(y^{L},y^{U},x;C)$ in Equation (12), let $s_{j}=s(Y^{L}_{j},Y^{U}_{j},X_{j};\hat{C})$, and $s_{j}^{*}=s\left(Y_{j}^{L},Y_{j}^{U},X_{j};C^{*}\right)$.
@@
Given a consistent estimator $\hat{C}$ for the oracle prediction set, if $s^{*}=s\left(Y^{L},Y^{U},X;C^{*}\right)$ has positive density in a neighbourhood around 00, then the conformal prediction set $\tilde{C}$ constructed as above, for an i.i.d. sample $\mathcal{D}=\left\{\left(X_{i},Y_{i}^{L},Y^{U}_{i}\right):i\in\mathcal{I}\right\}$ which is split into training $\mathcal{I}_{1}$ and calibration sets $\mathcal{I}_{2}$, we have
@@
as $n_{1},n_{2}\to\infty$, here $n_{1}=\left\lvert\mathcal{I}_{1}\right\rvert$ and $n_{2}=\left\lvert\mathcal{I}_{2}\right\rvert$.
@@
However, it is impossible to construct a prediction set that is finite-sample conditionally valid and consistent for the oracle prediction set (Lei and Wasserman,, 2014; Barber et al.,, 2021; Gibbs et al.,, 2024), and hence we consider the weaker notion of local validity. Let $\mathcal{A}=\left\{A_{k}:k\leq K\right\}$ be a partition of $\mathcal{X}$, a prediction interval ${C}_{\text{loc}}(x)$ is finite-sample locally valid with respect to $\mathcal{A}$, if for all $P$,
@@
$\vartheta_{1-\alpha,k}=q\left(\left\{s_{j}:j\in\mathcal{I}^{\mathcal{A}}_{k}\right\};(1-\alpha)\left(1+\frac{1}{n_{2,k}}\right)\right),$
@@
$\tilde{C}_{\text{loc}}(x)=\bigcup_{m=1}^{M}\left[\hat{\tau}_{0m}(x)-\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k},\hat{\tau}_{1m}(x)+\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k}\right].$
The finite-sample local validity is a result of exchangeability, similar to Lei and Wasserman, (2014).
@@ 5 Numerical experiments
  • Model A. The unobserved outcomes $Y_{i}$ are generated according to a mixture model with heteroscedasticity, similar to the simulations in Lei and Wasserman, (2014).
@@
    In this design, the conditional distribution $P_{Y\mid X}$ is unimodal for smaller values of $X$ and bi-modal for larger values of $X$. Regression and quantile regression based prediction sets will have difficulty capturing the bifurcation shape of the oracle prediction set.
  • Model B. The unobserved outcomes are generated in the same way as in Equation 14 above. We randomly select 20% of the unobserved outcomes and they are interval valued and in this case and $(Y^{L}_{i},Y_{i}^{U})$ have a discrete distribution , while the remaining 80% are observed without censoring. Let $\iota_{i}$ be a random draw from a Bernoulli distribution with probability 0.2,
@@
We then compare the average coverage and volume of the conformal prediction set $\tilde{C}$ and the local conformal prediction set $\tilde{C}_{\text{loc}}$ over 100 repetitions with two alternative conformal prediction intervals: $\tilde{C}_{q2}$, based on a quantile regression assuming a quadratic model, and $\tilde{C}_{q3}$, based on a quantile regression assuming a cubic model. See Romano et al., (2019) and Appendix E for more details about construction of prediction interval based on quantile methods. The coverage is computed by randomly generating 5,000 samples of $(Y^{L},Y^{U})$ and calculating the proportion of intervals covered by the prediction sets. The volume is calculated as the integrated Lebesgue measure of the prediction sets over the support $\mathcal{X}$.
Figure 3 shows the comparison of the prediction sets from the numerical experiments. The prediction sets $\tilde{C}$ and $\tilde{C}_{\text{loc}}$ capture the bifurcation region of the conditional distribution in Model A and Model B and have smaller volume compared to the quantile regression-based prediction sets. Even when the conditional distribution is uni-modal, and the oracle prediction set is a single interval for each $x\in\mathcal{X}$, our proposed estimators still have smaller volume as demonstrated in Model C. In terms of coverage, $\tilde{C},\tilde{C}_{\text{loc}}$ and $\tilde{C}_{q3}$ appear to have coverage close to the nominal level $0.91-\alpha=$ 0.9. Notably, $\tilde{C}_{\text{loc}}$ appears to have better coverage than $\tilde{C}$ near the boundaries of the support $\mathcal{X}$.
@@ 6.1 US CPS data
We applied our method to the 2023 US Current Population Survey (CPS) dataset and construct prediction interval for their income given years of education and experience. Following Heckman et al., (2008), we construct the following variables from the dataset for individuals who report being American citizen and employed:
  • Income: We include all individuals who reported their income or income range in the survey. Since 2014, when a person refuses to report their income, the interviewer will ask for a range, and the person’s income will be imputed by matching with similar people that have income in the same range. In the 2019 ASEC Updates [^4] See https://cps.ipums.org/cps/asec_2019_changes.shtml, it is stated that “if respondents did not give a value for income received from a given source, the interviewer now follows up with a question about income ranges. Respondents who did not give a specific value were asked if they received over \$60,000, between \$45,000 and \$60,000, or less than \$45,000 from this income source. Those who replied that they earned less than \$45,000 from this source were further asked if they earned more than \$30,000, between \$15,000 and \$30,000, or less than \$15,000 from this income source.” For people who reported their income as larger than \$$\$60,000$, we set the upper bound of the income as \$$\$400,000$, which is the topcode of the income variable in the dataset. The individuals who did not report their income nor their income range are excluded.
@@
  • Education: Following Heckman et al., (2008), we construct the years of education for each individual by converting the highest qualification they have obtained into years of education.
@@
The CPS dataset consists of 146,133 observations. There are 53,171 individuals who satisfy the demographic restrictions above. Among the individuals that satisfy the demographic criteria, 39,333 reported their income either exactly or as a band and the response rate is thus around 74%. The issue of nonresponse in surveys has been studied in the literature; see for example, Lillard et al., (1986) and Bollinger et al., (2019), where they document an increasing proportion of nonresponse of earnings in survey data and the potential bias it introduces. We will not address the nonresponse issue in this paper, but it is an important consideration in practice. Our proposed predictions could assist in the imputation of missing values in a robust manner; see Lei and Candès, (2021). In fact, allowing respondents to report their income in bands could help mitigate the non-response problem. For the 39,333 individuals who reported their incomes and satisfy the demographic restrictions, 8,819 of them reported either their gross or net income as a band, which thus amounts to roughly 22.4%. of the respondents. Figure 2 shows the proportion of individuals who reported their income in a range by years of education.
@@
In our empirical study, we first randomly reserve 20% of the data as a hold-out dataset. The remaining 80% is then further divided into training and calibration sets, with 75% of the data used for training to estimate the prediction interval $\hat{C}_{I}$ and 25% used for constructing the conformalized prediction interval $\tilde{C}_{I}$. Table 1 presents the average estimated prediction intervals across the 100 iterations. The method column indicates the way we estimate the prediction interval. “P” stands for prediction interval $\hat{C}_{I}$ in (7) and “CP” stands for conformal prediction interval $\tilde{C}_{I}$. If “M” is appended, it indicates the prediction intervals are estimated with imputed income for individuals who reported their income in an interval. It can be seen from the table that we would in general predict a higher income for individuals with more years of education.
@@
Table 2 shows the coverage of the interval valued incomes in the hold-out dataset for the conformal prediction interval $\tilde{C}$ that are constructed using both interval-valued and exactly reported income data (shown in the “Censored” columns), as well as for the conformal prediction intervals constructed when the interval-valued incomes are imputed (shown in the “Imputed” columns). The significant difference in prediction intervals constructed with interval-valued data and with imputed data may be due to the fact that there are only five categories for individuals who report their income in a range. This seems somewhat coarse in terms of informing about the true income distribution. Notice that the conformal coverage using the censored model has coverage close to the nominal rate while the Imputed Coverage is almost always lower. Note that in Tables 1 an 2 when we display different level of coverage $\alpha$, this provides an empirical estimate of the $(1-\alpha)$prediction sets which can be informative about the shape of the data distribution. Notice that standard reporting of confidence intervals on parameters is meant to summarize sampling uncertainty, while here not only do these conformalized sets take into account sampling uncertainty, but they also estimate prediction sets of the underlying distributions of interest.
7 Conclusion
The object of interest in this paper is a prediction set for an outcome given a set of covariates when this outcome is censored. Censoring is a common issue in economic and other data. We first characterise the oracle prediction set which is the smallest set that maintains a given miscoverage level under censoring. The characterisation leads to a feasible estimation strategy based on the observed sample. We then provide consistent estimators for this oracle prediction set based nonparametric estimation of the conditional distribution of the observed lower and upper brackets given the conditioning variables. We allow for the prediction set to consist of multiple intervals and consider both random and fixed censoring. Furthermore, we use recent results from conformal inference to obtain a conformal prediction set that maintains a finite sample miscoverage property using a set based score function. We find that our procedures perform well with simulated data. We also apply our procedures to data from the US Census. There is increasing interest in prediction set estimation, and several intriguing research directions are related to this paper. One promising avenue is to consider prediction sets under partial identification beyond interval censoring. While conformal inference offers finite-sample guarantees on coverage, it remains a research challenge to explore the more traditional statistical inference on the estimated prediction set. Additionally, the design of conformity scores in the context of partial identification presents an interesting and important topic for further investigation.
Model A Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model B Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model C Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption

### v4

@@ 4 Conformal inference with interval outcomes
where $\operatorname{P}$ denotes the joint probability of $\left\{\bar{Z}_{i}=(X_{i},Y_{i},Y^{L}_{i},Y_{i}^{U}):i\in\mathcal{I}\cup\left\{{n+1}\right\}\right\}$. Notice that this is the finite-sample validity condition typically considered in conformal inference literature, where the probability is the joint distribution for both the observed sample and the individual to be predicted. While the oracle prediction set $C^{*}(x)$ depends on the true distribution $P$, conformal inference constructions lead to a prediction set that is finite-sample valid for all $P$, making it distribution-free.
In particular, we will achieve finite-sample validity using the method of split conformal inference (see Romano, Patterson, and Candes (2019) and Lei, G’Sell, Rinaldo, Tibshirani, and Wasserman (2018)). One feasible split conformal inference procedure for the oracle prediction interval $C^{*}_{I}$ in our setting can be described as follows. We first split the sample $\mathcal{D}$ into a training set $\mathcal{D}_{1}=\left\{Z_{i}:i\in\mathcal{I}_{1}\right\}$ and a calibration set $\mathcal{D}_{2}=\left\{Z_{i}:i\in\mathcal{I}_{2}\right\}$. We calculate the estimator $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$ for the oracle prediction set with the training set for $x\in\mathcal{X}$ as in Equation 7. The following scores $s_{j}=s(Y^{L}_{j},Y^{U}_{j},X_{j};\hat{\tau}_{0},\hat{\tau}_{1})$ are then computed for each $j\in\mathcal{I}_{2}$ in the calibration set, given $\hat{C}_{I}(x)=[\hat{\tau}_{0}(x),\hat{\tau}_{1}(x)]$, where
@@
This score function is a modified version of the score function used in the case of conformalized quantile regression with no interval censoring studied in Romano et al. (2019). The score function is negative when the interval $y^{L},y^{U}]$ is contained in $t_{0}(x),t_{1}(x)]$ and positive otherwise. By allowing for negative scores, the conformal prediction set $\tilde{C}_{I}(x)$ penalises both overcoverage and undercoverage.
@@
[…]
@@
$s(y^{L},y^{U},x;C)=\inf\left\{\theta:[y^{L},y^{U}]\subset\bigcup_{m=1}^{M}[t_{0m}(x)-\theta,t_{1m}(x)+\theta]\right\}.$ (12)
This score can be computed explicitly as follows. With the notation $t_{10}=-\infty,t_{0,M+1}=\infty$, if there exists $j:1\leq j\leq M$ such that $t_{0,j}\leq y^{L}\leq y^{U}\leq t_{1,j}$, then $s(y^{L},y^{U};C)=s(y^{L},y^{U};t_{0j},t_{1j})$, otherwise, for $i=\max\{j:t_{1,j-1}\leq y^{L}\}$, and $k=\min\{j:t_{0,j+1}\geq y^{u}\}$,
$s(y^{L},y^{U},x;C)=$ $2^{-1}\max\{(t_{0,j}-t_{1,j-1}):[t_{1,j-1},t_{0,j}]\subset[y^{L},y^{U}]\}$
$\vee[(t_{0,i}-y^{L})\wedge 2^{-1}(t_{0,i}-t_{1,i-1})]$
$\vee[(y^{U}-t_{1,k})\wedge 2^{-1}(t_{0,k+1}-t_{1,k})].$
Notice that $s_{j}\leq 0$ if and only if $Y^{L}_{j},Y^{U}_{j}]\subset[t_{0m},t_{1m}]$ for some $m\leq M$. We can then compute $)\vartheta_{1-\alpha}=q(\{s_{j}:j\in\mathcal{I}_{2}\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert$)). The conformal prediction set is then given by  [^2] Here, the endpoints of the estimated $\hat{C}$ are adjusted uniformly using the same $\vartheta_{1-\alpha}$. A more sophisticated adjustment method is discussed in Appendix C, where we propose a minimal Euclidean distance adjustment. We conjecture that this approach could improve the convergence rate of the conformal prediction set. The optimal choice of conformity score remains an interesting question, see for example, Thurin et al. (2025).
@@
Under the assumption that $\{(X_{i},Y_{i},Y_{i}^{L},Y_{i}^{U}):i\in\mathcal{I}\cup\{n+1\}\}$ are i.i.d. [^3] In order to perform conformal inference, it is sufficient that the sample satisfies the weaker condition of exchangeability. Here, independence is assumed for simplicity and for establishing the asymptotics for the nonparametric estimators we will use. Barber et al. (2023) studies conformal inference beyond exchangeability and is applicable for time series data. Our results can be readily extended to the time series settings combined with their procedure. , the conformal prediction set $\tilde{C}$ in Equation 13 is finite-sample marginally valid, that is, for all $P$,
@@
[…]
@@
Given an i.i.d. sample $\mathcal{D}=\{(X_{i},Y_{i}^{L},Y^{U}_{i}):i\in\mathcal{I}\}$ which is split into training $\mathcal{I}_{1}$ and calibration sets $\mathcal{I}_{2}$. Let $\tilde{C}(x)$ be the conformal prediction set constructed by the preceding procedure in Equation 13, if $s(Y^{L},Y^{U},X;C^{*})$ has positive density in a neighbourhood around 0, then under the assumptions of Theorem 3.3,
@@
as the sample sizes $n_{1},n_{2}\to\infty$, here $n_{1}=\left\lvert\mathcal{I}_{1}\right\rvert$ and $n_{2}=\left\lvert\mathcal{I}_{2}\right\rvert$.
@@
However, it is impossible to construct a prediction set that is finite-sample conditionally valid and consistent for the oracle prediction set (Lei and Wasserman, 2014; Barber et al., 2021; Gibbs et al., 2024), and hence we consider the weaker notion of local validity. Let $\mathcal{A}=\left\{A_{k}:k\leq K\right\}$ be a partition of $\mathcal{X}$, a prediction interval ${C}_{\text{loc}}(x)$ is finite-sample locally valid with respect to $\mathcal{A}$, if for all $P$,
@@
$\vartheta_{1-\alpha,k}=q\left(\left\{s_{j}:j\in\mathcal{I}_{2,k}\right\};(1-\alpha)\left(1+\frac{1}{n_{2,k}}\right)\right),$
@@
$\tilde{C}_{\text{loc}}(x)=\bigcup_{m=1}^{M}\left[\hat{\tau}_{0m}(x)-\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k},\hat{\tau}_{1m}(x)+\sum_{k\leq K}\mathbb{I}\left\{x\in A_{k}\right\}\vartheta_{1-\alpha,k}\right].$ (14)
The finite-sample local validity is a result of exchangeability, similar to Lei and Wasserman (2014).
@@ 5 Numerical experiments
  • Model A. The unobserved outcomes $Y_{i}$ are generated according to a mixture model with heteroscedasticity, similar to the simulations in Lei and Wasserman (2014).
@@
    In this design, the conditional distribution $P_{Y\mid X}$ is unimodal for smaller values of $X$ and bi-modal for larger values of $X$. Regression and quantile regression-based prediction sets will have difficulty capturing the bifurcation shape of the oracle prediction set.
  • Model B. The unobserved outcomes are generated in the same way as in Equation 15 above. We randomly select 20% of the unobserved outcomes, and they are interval valued, and in this case and $(Y^{L}_{i},Y_{i}^{U})$ have a discrete distribution, while the remaining 80% are observed without censoring. Let $\iota_{i}$ be a random draw from a Bernoulli distribution with probability 0.2,
@@
We then compare the average coverage and volume of the conformal prediction set $\tilde{C}$ and the local conformal prediction set $\tilde{C}_{\text{loc}}$ over 100 repetitions with two alternative conformal prediction intervals: $\tilde{C}_{q2}$, based on a quantile regression assuming a quadratic model, and $\tilde{C}_{q3}$, based on a quantile regression assuming a cubic model. See Romano et al. (2019) and Appendix E for more details about the construction of a prediction interval based on quantile methods. The coverage is computed by randomly generating 5,000 samples of $(Y^{L},Y^{U})$ and calculating the proportion of intervals covered by the prediction sets. The volume is calculated as the integrated Lebesgue measure of the prediction sets over the support $\mathcal{X}$.
Figure 2 shows the comparison of the prediction sets from the numerical experiments. The prediction sets $\tilde{C}$ and $\tilde{C}_{\text{loc}}$ capture the bifurcation region of the conditional distribution in Model A and Model B and have a smaller volume compared to the quantile regression-based prediction sets. Even when the conditional distribution is unimodal, and the oracle prediction set is a single interval for each $x\in\mathcal{X}$, our proposed estimators still have smaller volume as demonstrated in Model C. In terms of coverage, $\tilde{C},\tilde{C}_{\text{loc}}$ and $\tilde{C}_{q3}$ appear to have coverage close to the nominal level $1-\alpha=0.9$. Notably, $\tilde{C}_{\text{loc}}$ appears to have better coverage than $\tilde{C}$ near the boundaries of the support $\mathcal{X}$.
Model A Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Model B Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Model C Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 6 Empirical study
6.1 UK Job Advert Data
We have also applied our method to the UK job advert dataset from Adzuna, an online job advert aggregator in the UK. Adzuna data have been used widely as an indicator of economic activity in the UK, including by the Office for National Statistics as experimental statistics (ONS, 2021). It has also been used to develop a measure of labour market opportunities for heterogeneous types of workers (Dias et al., 2021).
The original sample consists of job adverts posted in the United Kingdom between January and March of 2022, which comprise 121,746 observations. Each advertisement contains information on job location, job category, and salary, when available. Salary information is missing for 37.8% (46,077) of the postings. Among the advertisements that report salaries, the majority are listed as salary intervals, accounting for approximately 62% (46,606) of all postings with salary information, making this dataset well-suited for our interval-valued outcome framework.
The outcome of interest is the annual salary associated with each job advert, with $(Y_{i}^{L},Y_{i}^{U})$ denoting the reported lower and upper salary bounds for the $i$-th advert, and we work with the logarithm of the salary bounds. In the analysis below, we focus on the IT industry, which is the largest job category by number of postings (see Figure 3). For the 19,385 job advertisements in the IT industry, 40% (7,742 postings) report salary information, of which 64% (4,981 postings) provide salaries in intervals; in our analysis, we restrict attention to the subset of postings that report salary information. Job location is represented by geographic coordinates (latitude and longitude), which we treat as continuous covariates.
For a given location (latitude and longitude), we estimate the 90% oracle prediction interval (see 2.2) for the salary distribution in the IT category conditional on location. We employ the kernel smoothing estimator of the conditional probability using a rectangular kernel applied to geographic distance and solve for the empirical analogue of the feasible optimization problem in Equation (6), yielding prediction interval $\hat{C}=(\hat{\tau}_{0}(x),\hat{\tau}_{1}(x))$. Then we construct the local conformal prediction set $\tilde{C}_{\text{local}}$ according to Equation 14. To visualise spatial variation in predicted salary distributions, we evaluate the predicted upper and lower bounds on a grid of locations covering the UK and the resulting prediction bounds are then interpolated. Importantly, our approach accommodates the interval-valued nature of the salary data without imposing parametric assumptions or relying on imputation of missing salaries.
Figure 4 plots the estimated lower and upper bounds of the 90% local conformal prediction interval for annual salaries across geographic locations in the UK. For reference, the median annual pay for full-time employees was £33,000 for the tax year ending on 5 April 2022.444Source: https://www.ons.gov.uk/employmentandlabourmarket/peopleinwork/earningsandworkinghours/bulletins/annualsurveyofhoursandearnings/2022. Employee earnings in the UK: 2022, Measures of employee earnings using data from the Annual Survey of Hours and Earnings (ASHE), Office for National Statistics. We can observe a few patterns in the spatial distribution of the predicted salary bounds. First, the estimated lower bounds exhibit moderate geographic heterogeneity and are relatively even across space and are typically below £30,000. Lower bounds tend to be higher around major cities, while rural areas display lower predicted lower bounds or no job postings. Second, the upper bounds display a much thicker right tail and substantially stronger spatial concentration, with predicted values ranging roughly from £50,000 to £230,000. Areas surrounding major metropolitan centres, such as London, Manchester, Bristol, Birmingham and Edinburgh, stand out with markedly higher upper bounds than the rest of the country, indicating pronounced geographic concentration of high-paying jobs. As a result, the predicted salary interval is widest in London and other large cities, reflecting greater wage dispersion and job heterogeneity in these labour markets.
Table 1 reports the estimated 90% local conformal prediction intervals for annual IT salaries at selected major UK cities, evaluated at their geographic coordinates. The results illustrate substantial cross-city heterogeneity, particularly in the upper bounds of the predicted salary distributions. London stands out with both relatively higher predicted lower bound and upper bound, leading to the widest prediction interval among the listed cities, consistent with high wage dispersion and job heterogeneity in the capital. Other large metropolitan areas such as Manchester, Edinburgh, and Glasgow also exhibit wide intervals driven primarily by elevated upper bounds, while Birmingham displays a comparatively narrower interval, reflecting more moderate dispersion in predicted salaries. Across all cities, predicted lower bounds are relatively similar and generally below £25,000, whereas differences in upper bounds account for most of the variation in interval widths. We also note that, although the salary intervals reported in individual job postings are typically narrow, their levels vary substantially across postings even within the IT industry. This dispersion leads to the relatively large estimated interval widths.
Overall, these findings highlight the importance of accounting for interval-valued salary information when studying spatial variation in labour market opportunities. In particular, these patterns reinforce the importance of analysing the full conditional salary distribution rather than relying on point estimates, as the upper tail of the distribution varies markedly across locations. Our approach captures spatial variation in both the predicted lower and upper bounds, thereby providing richer insights than simple point estimates such as conditional means or medians, without imposing parametric assumptions or relying on imputation of missing salaries.
Table 1: Estimated 90% Local Conformal Prediction Intervals for IT Salaries in Major UK Cities
City Latitude Longitude Predicted lower Predicted upper Interval width
London 51.50°N 0.14°W £22,204 £180,094 £157,889
Manchester 53.48°N 2.23°W £19,757 £159,553 £139,795
Birmingham 52.48°N 1.89°W £16,885 £110,028 £93,143
Edinburgh 55.94°N 3.33°W £17,855 £159,388 £141,533
Glasgow 55.86°N 4.26°W £15,822 £144,000 £128,178
Bristol 51.46°N 2.59°W £22,636 £135,657 £113,021
Figure 3: Job advert counts by category.
(a) Lower bound
(b) Upper bound
Figure 4: Local conformal prediction interval for annual salaries in the UK IT job market. The left panel shows the estimated lower bound, and the right panel shows the estimated upper bound of the 90% local conformal prediction interval for annual salaries across geographic locations in the UK
@@ 6.2 US CPS data
We apply our method to the 2023 US Current Population Survey (CPS) dataset and construct a prediction interval for their income given years of education and experience. Following Heckman et al. (2008), we construct the following variables from the dataset for individuals who report being American citizens and employed:
  • Income: We include all individuals who reported their income or income range in the survey. Since 2014, when a person refuses to report their income, the interviewer will ask for a range, and the person’s income will be imputed by matching with similar people who have income in the same range. In the 2019 ASEC Updates [^5] See https://cps.ipums.org/cps/asec_2019_changes.shtml, it is stated that “if respondents did not give a value for income received from a given source, the interviewer now follows up with a question about income ranges. Respondents who did not give a specific value were asked if they received over \$60,000, between \$45,000 and \$60,000, or less than \$45,000 from this income source. Those who replied that they earned less than \$45,000 from this source were further asked if they earned more than \$30,000, between \$15,000 and \$30,000, or less than \$15,000 from this income source.” For people who reported their income as larger than $\mathdollar 60,000$, we set the upper bound of the income as $\mathdollar 400,000$, which is the topcode of the income variable in the dataset. The individuals who did not report their income or their income range are excluded.
@@
  • Education: Following Heckman et al. (2008), we construct the years of education for each individual by converting the highest qualification they have obtained into years of education.
@@
The CPS dataset consists of 146,133 observations. There are 53,171 individuals who satisfy the demographic restrictions above. Among the individuals that satisfy the demographic criteria, 39,333 reported their income either exactly or as a band and the response rate is thus around 74%. The issue of nonresponse in surveys has been studied in the literature; see for example, Lillard et al. (1986) and Bollinger et al. (2019), where they document an increasing proportion of nonresponse of earnings in survey data and the potential bias it introduces. We will not address the nonresponse issue in this paper, but it is an important consideration in practice. Our proposed predictions could assist in the imputation of missing values in a robust manner; see Lei and Candès (2021). In fact, allowing respondents to report their income in bands could help mitigate the non-response problem. For the 39,333 individuals who reported their incomes and satisfy the demographic restrictions, 8,819 of them reported either their gross or net income as a band, which thus amounts to roughly 22.4%. of the respondents. Figure 5 shows the proportion of individuals who reported their income in a range by years of education.
@@
In our empirical study, we first randomly reserve 20% of the data as a hold-out dataset. The remaining 80% is then further divided into training and calibration sets, with 75% of the data used for training to estimate the prediction interval $\hat{C}_{I}$ and 25% used for constructing the conformalized prediction interval $\tilde{C}_{I}$. Table 2 presents the average estimated prediction intervals across the 100 iterations. The method column indicates the way we estimate the prediction interval. “P” stands for prediction interval $\hat{C}_{I}$ in (7) and “CP” stands for conformal prediction interval $\tilde{C}_{I}$. If “M” is appended, it indicates the prediction intervals are estimated with imputed income for individuals who reported their income in an interval. It can be seen from the table that we would in general, predict a higher income for individuals with more years of education.
@@
Table 3 shows the coverage of the interval valued incomes in the hold-out dataset for the conformal prediction interval $\tilde{C}$ that are constructed using both interval-valued and exactly reported income data (shown in the “Censored” columns), as well as for the conformal prediction intervals constructed when the interval-valued incomes are imputed (shown in the “Imputed” columns). The significant difference in prediction intervals constructed with interval-valued data and with imputed data may be because there are only five categories for individuals who report their income in a range. This seems somewhat coarse in terms of informing about the true income distribution. Notice that the conformal coverage using the censored model has coverage close to the nominal rate, while the Imputed Coverage is almost always lower. Note that in Tables 2 and 3, when we display different levels of coverage $\alpha$, this provides an empirical estimate of the $(1-\alpha)$% prediction sets, which can be informative about the shape of the data distribution. Notice that standard reporting of confidence intervals on parameters is meant to summarise sampling uncertainty, while here not only do these conformalized sets take into account sampling uncertainty, but they also estimate prediction sets of the underlying distributions of interest.